OpenStudy (anonymous):

Find a quadratic function that has zeroes -4 and 9 and a graph that passes through (3,5).

OpenStudy (anonymous):

y = (x + 0) (x+4) (x-9) <-do you see how I set this up given the info you had? y = x^3 - 5x^2 - 36x <- just simplifying it out now, adapt the equation to meet the point (3,5) y = (3)^3 - 5 (3)^2 - 36 (3) y = -126 so we need y to be 5, not -126, so we will add 131 y = x^3 - 5x^2 - 36x +131 y = (3)^3 - 5 (3)^2 - 36 (3) +131 y = 5 make sense?

OpenStudy (anonymous):

How did you get x+0? 0 isn't a root and it says /quadratic equation/

OpenStudy (anonymous):

I'm probably misunderstanding the question then, i though it said zeros

OpenStudy (anonymous):

zeroes meaning roots. But I've finished the problem now.

OpenStudy (anonymous):

ok! ^_^